# PIPER: Primitive-Informed Preference-based Hierarchical Reinforcement Learning via Hindsight Relabeling

### v1

@@ 1 Introduction
Off-policy hierarchical reinforcement learning (HRL) (Sutton et al., 1999; Dayan & Hinton, 1992; Vezhnevets et al., 2017; Klissarov et al., 2017; Harb et al., 2018) approaches hold the promise of improved sample efficiency due to temporal abstraction and improved performance due to more effective exploration (Nachum et al., 2019). In goal-conditioned feudal architectures (Dayan & Hinton, 1992; Vezhnevets et al., 2017), a widely used subset of HRL methods, higher-level policies predict subgoals for the lower-level policies, while the lower-level policies try to achieve those subgoals by executing primitive actions directly in the environment. Despite their advantages, serious challenges remain for such approaches. In this paper, we focus on two important issues: the destabilizing effect of lower-level reward non-stationarity on off-policy HRL (Levy et al., 2018; Nachum et al., 2018b), and performance degradation due to infeasible subgoal generation by higher-level policies Chane-Sane et al. (2021).
Recent approaches in preference-based learning Christiano et al. (2017); Ibarz et al. (2018); Lee et al. (2021) use human feedback to learn a reward model and subsequently apply RL to solve the task using the learned reward model. Impressively, such methods have demonstrated performance comparable to RL policies trained with access to ground truth rewards. This raises the following question: can advances in preference-based learning be used to simultaneously address the twin issues of reward non-stationarity and infeasible subgoal generation in HRL?
@@
Major contributions of PIPER:
@@ 2 Related Work
The option learning framework (Sutton et al., 1999; Klissarov et al., 2017) provides the benefits of temporal abstraction by learning extended macro actions. Unfortunately, such approaches can lead to degenerate solutions and thus require additional regularization approaches to ensure feasible subgoal prediction. Notably, such degenerate solutions result in unbalanced task-split which nullifies the advantages of hierarchical learning. In our approach, we perform primitive-informed regularization to regularize the higher-level policy to predict achievable subgoals for lower primitive. Another line of work uses previously hand-designed action primitives Nasiriany et al. (2021); Dalal et al. (2021) to accelerate hierarchical learning. Such approaches, however, depend on the quality of hand-designed primitives, which is difficult to design in complex environments.
@@ 3.1 Limitations of Existing Approaches to HRL
While HRL promises significant advantages over non-hierarchical RL, including improved sample efficiency due to temporal abstraction and improved exploration Nachum et al. (2018b, 2019), serious challenges remain. In this work, we focus on two outstanding issues:
@@ 4.1 Preference-based Learning (PBL)
In traditional RL, agents learn to maximize the accumulated rewards $r(s_{t},a_{t})$ obtained from the environment, where the reward $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is assumed to be known. In many real-world scenarios, however, suitable environment rewards are notoriously difficult to construct and typically require domain-specific knowledge Ng et al. (1999). To learn the high-level reward necessary for the HRL setting considered in this paper, we use a preference-based RL setup, where the agent learns to perform the high-level task using preferences over agent behaviors (Wilson et al., 2012a; Christiano et al., 2017; Lee et al., 2021; Ibarz et al., 2018).
@@
where $\delta^{1}\succ\delta^{2}$ denotes the event that $\delta^{1}$ is preferred over $\delta^{2}$. To learn parameters $\phi$ such that (2) matches the true preferences, behavior and preference data are recorded in a dataset $\mathcal{D}$ with entries of the form $(\delta^{1},\delta^{2},y)$, where $y=(1,0)$ when $\delta^{1}$ is preferred over $\delta^{2}$, $y=(0,1)$ when $\delta^{2}$ is preferred over $\delta^{1}$, and $y=(0.5,0.5)$ when there is no preference. The standard approach in the preference-based literature (see Christiano et al. (2017); Lee et al. (2021)), which we adopt in PIPER, is to learn the reward function $\widehat{r}_{\phi}$ using a cross-entropy loss:
$\small\operatorname{L}(\phi)=-\sum_{\mathcal{D}}\left(y_{1}\log P_{\phi}\left[\delta^{1}\succ\delta^{2}\right]+y_{2}\log P_{\phi}\left[\delta^{2}\succ\delta^{1}\right]\right),$ (3)
@@
  • Finally, we incorporate soft target updates Lillicrap et al. (2015) to mitigate the training instability while learning the high-level reward function. Taken together, this combination of techniques constitutes PIPER, the primary contribution of this paper. The rest of this section provides details on each aspect of our approach. Pseudocode for PIPER is provided in Algorithm 1.
@@ 4.2 PiL: Primitive-in-the-Loop Feedback for HRL
$\small\sum_{i=0}^{n-1}$ $r(s^{1}_{t+ik},g^{*},g^{1}_{t+ik})>\sum_{i=0}^{n-1}r(s^{2}_{t+ik},g^{*},g^{2}_{t+ik})$ (4)
Replacing human feedback: In the standard preference-learning framework, preferences are elicited from human feedback Christiano et al. (2017) and are subsequently used to learn the reward model $\widehat{r}_{\phi}$. In this work, we replace this Human-in-the-Loop (HiL) feedback with Primitive-in-the-Loop (PiL) feedback by using implicit sparse rewards, $r^{s}(s_{t},g^{*},g_{t})$, defined presently, to determine preferences $y$ between behavior sequences $\sigma^{1}$ and $\sigma^{2}$. We call this feedback Primitive-in-the-Loop since we generate this feedback using primitive sparse rewards. We obtain these primitive rewards as follows. Suppose the higher level policy $\pi^{H}$ predicts subgoal $g_{t}\sim\pi^{H}(\cdot|s_{t},g^{*})$ for state $s_{t}$ and goal $g^{*}$. The lower-level primitive executes primitive actions according to its policy $\pi^{L}$ for $k$ timesteps and ends up in state $s_{t+k-1}$. We use the sparse reward provided by the environment at state $s_{t+k-1}$ as the implicit sparse reward, i.e., $r^{s}(s_{t},g^{*},g_{t})=-\mathbf{1}_{\{\|s_{t+k-1}-g^{*}\|_{2}>\epsilon\}}$. Note that this reward is directly available from the environment. We replace the implicit reward $r$ in Equation (4) with $r^{s}$, and thus use $r^{s}$ to obtain preferences between higher level behavior sequences. Concretely, for the goal $g^{*}$, $\sigma^{1}\succ\sigma^{2}$ implies
$\small\sum_{i=0}^{n-1}$ $r^{s}(s^{1}_{t+ik},g^{*},g^{1}_{t+ik})>\sum_{i=0}^{n-1}r^{s}(s^{2}_{t+ik},g^{*},g^{2}_{t+ik})$ (5)
4.3 PiL with Goal-conditioned Hindsight Relabeling
4.4 Primitive-informed Regularization
For a given lower-level policy $\pi^{L}:\mathcal{S}\times\mathcal{G}\rightarrow\mathcal{A}$, denote the corresponding state value function by $V^{L}(s,g)$. To encourage appropriate subgoal selection by the higher-level policy, we propose using the lower-level state value function $V^{L}$ to regularize the preference reward developed in the previous section. Intuitively, $V^{L}(s_{t},g_{t})$ provides an estimate of the achievability of subgoal $g_{t}$ from current state $s_{t}$, since a high value of $V^{L}(s_{t},g_{t})$ implies that the lower level expects to achieve high reward in this setting. Concretely, we regularize the sparse reward $r^{s}$ defined in Section 4.2 with the function $r^{v}(s,g)=V^{L}(s,g)$ to obtain $r^{total}(s,g^{*},g)=r^{s}(s,g^{*},g)+\alpha r^{v}(s,g)$, where $\alpha\geq 0$ is a regularization hyperparameter. Substituting this primitive reward into inequality (6) yields our final preference condition for determining whether $\delta^{1}\succ\delta^{2}$:
@@ 4.4 Primitive-informed Regularization
10:     $\mathcal{R}^{H}\leftarrow\mathcal{R}^{H}\cup d^{H}$
11:     $\mathcal{R}^{L}\leftarrow\mathcal{R}^{L}\cup d^{L}$
12:     // Sample higher level behavior trajectories and goal
13:     $(\sigma^{1},\sigma^{2})\sim\mathcal{R^{H}}$ and $g\sim\mathcal{G}$
14:     Sample a set of additional goals for relabeling $\mathcal{G^{{}^{\prime}}}$
15:     for $g^{{}^{\prime}}\in\mathcal{G^{{}^{\prime}}}$ do
16:        Relabel $g$ by $g^{{}^{\prime}}$ and generate label $y$ using Equation (7)
17:        Store preference $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\sigma^{1},\sigma^{2},y)\}$
18:     // Reward Model Learning
19:     for each gradient step do
20:        Optimize model reward $\widehat{r}_{\phi}$ using Equation (3)
21:        // Soft update target reward model parameters
22:        $\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime}$
23:     // Policy Learning
24:     for each gradient step do
25:        Sample $\{(\sigma_{j})\}_{j=1}^{m}$ from $\mathcal{R}^{H}$
26:        Sample $\{(\delta_{j})\}_{j=1}^{m}$ from $\mathcal{R}^{L}$
@@
27:        Relabel rewards in $\{(\sigma_{j})\}_{j=1}^{m}$ using target reward model $\widehat{r}_{\phi^{{}^{\prime}}}$
28:        Optimize higher policy $\pi^{H}$ using SAC
29:        Optimize lower policy $\pi^{L}$ using SAC
@@ 4.5 PIPER Implementation
Pseudocode details: We explain our approach in detail in Algorithm 1. The higher- and lower-level policies are both trained using Soft Actor Critic (SAC) (Haarnoja et al., 2018). The rewards in higher level transitions are relabeled using the reward model $\widehat{r}_{\phi}$. Instead of relabeling all replay buffer transitions, we relabel the higher-level replay buffer transitions as they are sampled while training.
@@ 5 Experiments
(a) Maze navigation
(b) Pick and place
(c) Push
(d) Hollow
(e) Kitchen
@@ 5.1 Setup
We evaluate PIPER on four robotic navigation and manipulation tasks: $(i)$ maze navigation, $(ii)$ pick and place (Andrychowicz et al., 2017), $(iii)$ push, $(iv)$ hollow, and $(v)$ franka kitchen (Gupta et al., 2019). Since we focus on sparse reward scenarios, we re-implement these environments as sparse reward environments. For environments $(i)-(v)$, the maximum task horizon $\mathcal{T}$ is set to $225$, $50$, $50$, $100$, $225$ timesteps, respectively, and the lower primitive is allowed to execute for $k=15,7,7,10$ and $15$ timesteps, respectively. For environments $(i)-(v)$, the primitive regularization weight hyper-parameter $\alpha$ is set as $10^{-5}$, $10^{-4}$, $10^{-4}$, $10^{-4}$, $10^{-6}$ in environments $(i)-(v)$, respectively (see Figure 3). In our experiments, we use off-policy Soft Actor Critic (SAC) (Haarnoja et al., 2018) for optimizing RL objective, using the Adam (Kingma & Ba, 2014) optimizer. The actor and critic networks are formulated as three-layer, fully connected neural networks with $512$ neurons in each layer. The experiments are run for 6.75e5, 1.5e5, 6.5E5, and 6.75e5 timesteps in environments $(i)-(iv)$, respectively. Notably, since the pick and place, push, hollow and kitchen task environments are difficult, in order to speedup training, we assume access to a single human demonstration, and use an additional imitation learning objective at the lower level. We do not assume access to any demonstration in the maze navigation task. We keep this assumption consistent among all baselines to ascertain fair comparisons. We provide additional implementation details in Appendix A.2.

### v2

@@ 1 Introduction
Off-policy hierarchical reinforcement learning (HRL) (Sutton et al., 1999; Dayan & Hinton, 1992; Vezhnevets et al., 2017; Klissarov et al., 2017; Harb et al., 2018) approaches hold the promise of improved sample efficiency due to temporal abstraction and improved performance due to more effective exploration (Nachum et al., 2019). In goal-conditioned feudal architectures (Dayan & Hinton, 1992; Vezhnevets et al., 2017), a widely used subset of HRL methods, higher-level policies predict subgoals for the lower-level policies, while the lower-level policies try to achieve those subgoals by executing primitive actions directly in the environment. Despite their advantages, serious challenges remain for such approaches. In this paper, we focus on two important issues: the destabilizing effect of lower-level reward non-stationarity on off-policy HRL (Levy et al., 2018; Nachum et al., 2018b), and performance degradation due to infeasible subgoal generation by higher-level policies (Chane-Sane et al., 2021).  [^1] The implementation code and data is provided here.
Recent approaches in preference-based learning (Christiano et al., 2017; Ibarz et al., 2018; Lee et al., 2021) use human feedback to learn a reward model and subsequently apply RL to solve the task using the learned reward model. Impressively, such methods have demonstrated performance comparable to RL policies trained with access to ground truth rewards. This raises the following question: can advances in preference-based learning be used to simultaneously address the twin issues of reward non-stationarity and infeasible subgoal generation in HRL?
@@
We summarize the main contributions of PIPER:
@@ 2 Related Works
The option learning framework (Sutton et al., 1999; Klissarov et al., 2017) provides the benefits of temporal abstraction by learning extended macro actions. Unfortunately, such approaches can lead to degenerate solutions and thus require additional regularization approaches to ensure feasible subgoal prediction. Notably, such degenerate solutions result in unbalanced task-split which nullifies the advantages of hierarchical learning. In our approach, we perform primitive-informed regularization to regularize the higher-level policy to predict achievable subgoals for lower primitive. Another line of work uses previously hand-designed action primitives (Nasiriany et al., 2021; Dalal et al., 2021) to accelerate hierarchical learning. Such approaches, however, depend on the quality of hand-designed primitives, which is difficult to design in complex environments.
@@ 3.1 Limitations of Existing Approaches to HRL
While HRL promises significant advantages over non-hierarchical RL, including improved sample efficiency due to temporal abstraction and improved exploration (Nachum et al., 2018b, 2019), serious challenges remain. In this work, we focus on two outstanding issues:
@@ 4.1 Preference-based Learning (PBL)
In traditional RL, agents learn to maximize the accumulated rewards $r(s_{t},a_{t})$ obtained from the environment, where the reward $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is assumed to be known. In many real-world scenarios, however, suitable environment rewards are notoriously difficult to construct and typically require domain-specific knowledge (Ng et al., 1999). To learn the high-level reward necessary for the HRL setting considered in this paper, we use a preference-based RL setup, where the agent learns to perform the high-level task using preferences over agent behaviors (Wilson et al., 2012a; Christiano et al., 2017; Lee et al., 2021; Ibarz et al., 2018).
@@
where $\delta^{1}\succ\delta^{2}$ denotes the event that $\delta^{1}$ is preferred over $\delta^{2}$. To learn parameters $\phi$ such that (2) matches the true preferences, behavior and preference data are recorded in a dataset $\mathcal{D}$ with entries of the form $(\delta^{1},\delta^{2},y)$, where $y=(1,0)$ when $\delta^{1}$ is preferred over $\delta^{2}$, $y=(0,1)$ when $\delta^{2}$ is preferred over $\delta^{1}$, and $y=(0.5,0.5)$ when there is no preference. The standard approach in the preference-based literature (see (Christiano et al., 2017; Lee et al., 2021)), which we adopt in PIPER, is to learn the reward function $\widehat{r}_{\phi}$ using a cross-entropy loss:
$\operatorname{L}(\phi)=-\!\sum_{\mathcal{D}}\left(y_{1}\!\log P_{\phi}\left[\delta^{1}\!\succ\!\delta^{2}\right]+y_{2}\log P_{\phi}\left[\delta^{2}\!\succ\!\delta^{1}\right]\right),$ (3)
@@
  • Finally, we incorporate soft target updates (Lillicrap et al., 2015) to mitigate the training instability while learning the high-level reward function. Taken together, this combination of techniques constitutes PIPER, the primary contribution of this paper. The rest of this section provides details on each aspect of our approach. Pseudo-code for PIPER is provided in Algorithm 1.
@@ 4.2 PiL: Primitive-in-the-Loop Feedback for HRL
10:     end for
11:     $\mathcal{R}^{H}\leftarrow\mathcal{R}^{H}\cup d^{H}$
12:     $\mathcal{R}^{L}\leftarrow\mathcal{R}^{L}\cup d^{L}$
13:     // Sample higher level behavior trajectories and goal
14:     $(\sigma^{1},\sigma^{2})\sim\mathcal{R^{H}}$ and $g\sim\mathcal{G}$
15:     Sample a set of additional goals for relabeling $\mathcal{G^{{}^{\prime}}}$
16:     for $g^{{}^{\prime}}\in\mathcal{G^{{}^{\prime}}}$ do
17:        Relabel $g$ by $g^{{}^{\prime}}$ and generate label $y$ using Equation (11)
18:        Store preference $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\sigma^{1},\sigma^{2},y)\}$
19:     end for
20:     // Reward Model Learning
21:     for each gradient step do
22:        Optimize model reward $\widehat{r}_{\phi}$ using Equation (3)
23:        // Soft update target reward model parameters
24:        $\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime}$
25:     end for
26:     // Policy Learning
27:     for each gradient step do
28:        Sample $\{(\sigma_{j})\}_{j=1}^{m}$ from $\mathcal{R}^{H}$
29:        Sample $\{(\delta_{j})\}_{j=1}^{m}$ from $\mathcal{R}^{L}$
@@
30:        Relabel rewards in $\{(\sigma_{j})\}_{j=1}^{m}$ using target reward model $\widehat{r}_{\phi^{{}^{\prime}}}$
31:        Optimize higher policy $\pi^{H}$ using SAC
32:        Optimize lower policy $\pi^{L}$ using SAC
33:     end for
34:  end for
$\small\sum_{i=0}^{n-1}$ $r(s^{1}_{t+ik},g^{*},g^{1}_{t+ik})>\sum_{i=0}^{n-1}r(s^{2}_{t+ik},g^{*},g^{2}_{t+ik}).$ (4)
Replacing human feedback: In the standard preference-learning framework, preferences are elicited from human feedback (Christiano et al., 2017) and are subsequently used to learn the reward model $\widehat{r}_{\phi}$. In this work, we replace this Human-in-the-Loop (HiL) feedback with Primitive-in-the-Loop (PiL) feedback by using implicit sparse rewards, $r^{s}(s_{t},g^{*},g_{t})$, defined presently, to determine preferences $y$ between behavior sequences $\sigma^{1}$ and $\sigma^{2}$. We call this feedback Primitive-in-the-Loop since we generate this feedback using primitive sparse rewards. We obtain these primitive rewards as follows. Suppose the higher level policy $\pi^{H}$ predicts subgoal $g_{t}\sim\pi^{H}(\cdot|s_{t},g^{*})$ for state $s_{t}$ and goal $g^{*}$. The lower-level primitive executes primitive actions according to its policy $\pi^{L}$ for $k$ timesteps and ends up in state $s_{t+k-1}$. We use the sparse reward provided by the environment at state $s_{t+k-1}$ as the implicit sparse reward, i.e., $r^{s}(s_{t},g^{*},g_{t})=-\mathbf{1}_{\{\|s_{t+k-1}-g^{*}\|_{2}>\epsilon\}}$. Note that this reward is directly available from the environment. We replace the implicit reward $r$ in Equation (4) with $r^{s}$, and thus use $r^{s}$ to obtain preferences between higher level behavior sequences. Concretely, for the goal $g^{*}$, $\sigma^{1}\succ\sigma^{2}$ implies
$\small\sum_{i=0}^{n-1}$ $r^{s}(s^{1}_{t+ik},g^{*},g^{1}_{t+ik})>\sum_{i=0}^{n-1}r^{s}(s^{2}_{t+ik},g^{*},g^{2}_{t+ik}).$ (5)
4.3 PiL with Goal-conditioned Hindsight Relabeling
4.4 Primitive-informed Regularization
For a given lower-level policy $\pi^{L}:\mathcal{S}\times\mathcal{G}\rightarrow\mathcal{A}$, denote the corresponding state value function by $V_{\pi^{L}}(s,g)$. To encourage appropriate subgoal selection by the higher-level policy, we propose using the lower-level state value function $V_{\pi^{L}}$ to regularize the higher-level policy to predict feasible subgoals for the lower-level policy. Intuitively, $V_{\pi^{L}}(s_{t},g_{t})$ provides an estimate of the achievability of subgoal $g_{t}$ from current state $s_{t}$, since a high value of $V_{\pi^{L}}(s_{t},g_{t})$ implies that the lower level expects to achieve high reward for subgoal $g_{t}$. Let $r^{s}(s,g^{*},g)$ denote the parameterized reward model corresponding to the preference data, as defined in Section 4.2. For a trajectory $\tau$ of length $T$, consider the following KL-regularized formulation of preference-based learning, where $\pi_{reg}$ is the regularizing policy for the higher-level policy $\pi^{H}$:
$\max_{\pi^{H}}\mathbb{E}_{\pi^{H}}[\sum_{t=0}^{T}(r^{s}(s,g^{*},g)-\beta\mathbb{D}_{\mathrm{KL}}[\pi^{H}(\cdot|s_{t})\|\pi_{reg}(\cdot|s_{t})])],$ (7)
where $\beta\geq 0$ is a scalar hyperparameter controlling the deviation from the regularization policy $\pi_{reg}$. We propose the following formulation of the regularization policy:
$\pi_{{reg}}(g_{t}\mid s_{t})=\frac{\exp(m(V_{\pi^{L}}(s_{t},g_{t})))}{\hat{Z}(s_{t})},$ (8)
where $\hat{Z}(s_{t})=\sum_{g_{t}}\exp(m(V_{\pi^{L}}(s_{t},g_{t}))$, and $m=\frac{\alpha}{\beta}$. The policy $\pi_{reg}(\cdot|g_{t})$ is simply the softmax distribution generating a given subgoal $g_{t}$ with probability proportional to its value $V_{\pi^{L}}(s_{t},g_{t})$. We substitute (8) in (7) to get
$\begin{split}\max_{\pi^{H}}\mathbb{E}_{\pi^{H}}\left[\sum_{t=0}^{T}(r^{s}(s,g^{*},g)+\alpha(V_{\pi^{L}}(s_{t},g_{t}))+\hat{m}(s_{t}))\right],\end{split}$ (9)
where $\hat{m}(s_{t})=\beta\mathcal{H}(s_{t})-\beta\log Z(s_{t})$, and $\mathcal{H}(s_{t})=-\log\pi^{H}(g_{t}|s_{t})$ is the entropy term for $\pi^{H}$. Following prior work (Levine, 2018; Ziebart et al., 2008), we get the following optimal solution for the higher-level policy:
$\pi^{H}(g_{t}|s_{t})=\frac{1}{Z(s_{t})}\exp(\frac{1}{\beta}(r^{s}(s,g^{*},g)+\alpha(V_{\pi^{L}}(s_{t},g_{t})))),$ (10)
where $Z(s_{t})=\sum_{g_{t}}\exp(\frac{1}{\beta}(r^{s}(s,g^{*},g)+\alpha(V_{\pi^{L}}(s_{t},g_{t}))))$ is the partition function and $\alpha$ is the primitive regularization weight hyperparameter. Appendix A.1 contains the complete derivation. Notice that this optimal policy $\pi^{H}(g_{t}|s_{t})$ assigns high probability to subgoals $g_{t}$ which maximize the regularized reward $r^{total}(s,g^{*},g)=r^{s}(s,g^{*},g)+\alpha(V_{\pi^{L}}(s_{t},g_{t}))$. If we use $r^{total}$ to generate preferences between trajectories instead of the standard preferences outlined in Section 4.1, we end up with the optimal policy under this primitive-regularized, preference-based learning scheme. Hence, we substitute $r^{total}(s,g^{*},g)$ into inequality (6) to yield our final preference condition for determining if $\delta^{1}\succ\delta^{2}$:
@@ 4.5 PIPER Implementation
Pseudo-code details: We explain our approach in detail in Algorithm 1. The higher- and lower-level policies are both trained using Soft Actor Critic (SAC) (Haarnoja et al., 2018). The rewards in higher level transitions are relabeled using the reward model $\widehat{r}_{\phi}$. Instead of relabeling all replay buffer transitions, we relabel the higher-level replay buffer transitions as they are sampled while training.
@@ 5.1 Setup
We evaluate PIPER on four robotic navigation and manipulation tasks: $(i)$ maze navigation, $(ii)$ pick and place (Andrychowicz et al., 2017), $(iii)$ push, $(iv)$ hollow, and $(v)$ franka kitchen (Gupta et al., 2019). Since we focus on sparse reward scenarios, we re-implement these environments as sparse reward environments. Notably, since the pick and place, push, hollow and kitchen task environments are difficult, in order to speedup training, we assume access to a single human demonstration, and use an additional imitation learning objective at the lower level. We do not assume access to any demonstration in the maze navigation task. We keep this assumption consistent among all baselines to ascertain fair comparisons. We provide additional implementation details in Appendix A.3.
For environments $(i)-(v)$ ($(i)$ maze navigation, $(ii)$ pick and place (Andrychowicz et al., 2017), $(iii)$ push, $(iv)$ hollow, and $(v)$ franka kitchen (Gupta et al., 2019)), the maximum task horizon $\mathcal{T}$ is set to $225$, $50$, $50$, $100$, $225$ timesteps, respectively, and the lower primitive is allowed to execute for $k=15,7,7,10$ and $15$ timesteps, respectively. For environments $(i)-(v)$, the primitive regularization weight hyper-parameter $\alpha$ is set as $10^{-5}$, $10^{-4}$, $10^{-4}$, $10^{-4}$, $10^{-6}$ in environments $(i)-(v)$, respectively (see Figure 3). In our experiments, we use off-policy Soft Actor Critic (SAC) (Haarnoja et al., 2018) for optimizing RL objective, using the Adam (Kingma & Ba, 2014) optimizer. The actor and critic networks are formulated as three-layer, fully connected neural networks with $512$ neurons in each layer. The experiments are run for 6.75e5, 1.5e5, 6.5E5, and 6.75e5 timesteps in environments $(i)-(iv)$, respectively. In our experiments, we use off-policy Soft Actor Critic (SAC) (Haarnoja et al., 2018) for optimizing RL objective, using the Adam (Kingma & Ba, 2014) optimizer. The actor and critic networks are formulated as three-layer, fully connected neural networks with $512$ neurons in each layer. The experiments are run for 6.75e5, 1.5e5, 6.5E5, and 6.75e5 timesteps in environments $(i)-(iv)$, respectively.
In the maze navigation task, the closed gripper (fixed at table height) of a $7$-degree-of-freedom (7-DoF) robotic arm has to navigate across a four room maze to the goal position. In pick and place task, a $7$-DoF robotic arm gripper has to navigate to the square block, pick it up and bring it to the goal position. In the push task, a $7$-DoF robotic arm gripper has to push the square block towards the goal position. In hollow task, the $7$-DOF robotic arm gripper has to pick a square hollow block and place it such that a fixed vertical structure on the table goes through the hollow block. In the kitchen task, a $9$-DoF franka robot has to navigate and open the microwave door. To ensure fair comparisons, we keep parameters including neural network layer width, number of layers, choice of optimizer, SAC implementation parameters, etc., consistent across all baselines. For implementing the RAPS baseline, we use the following lower-level behaviors: in maze navigation, we design a single primitive, reach, where the lower-level primitive travels in a straight line towards the subgoal predicted by higher level. In the pick and place, push and hollow tasks, we design three primitives: gripper-reach, where the gripper goes to given position $(x_{i},y_{i},z_{i})$; gripper-open, which opens the gripper; and gripper-close, which closes the gripper. In the kitchen environment, we use the action primitives implemented in RAPS (Dalal et al., 2021).